OpenStudy (monicalabranche):

i am reviewing this for a college class i am taking . I am rusty and need some help and pointers ... x + 9 = 18 - 2x ​​

OpenStudy (utophii):

Add 2x - 9 to both side of the eqn. Then 3x = 9, dividing 3 both side, we obtain x = 3.

OpenStudy (mathwizzard3):

Let me know if you need help with anything like that :3

OpenStudy (monicalabranche):

Thanks

OpenStudy (laura.swart):

3x+9=18 3x=9 x=3

OpenStudy (macgirl234):

x + 9 = 18 - 2x 3x + 9 = 18 3x = 9 x = 3